OpenStudy (anonymous):

Anyone who can help would be amazing!! (-1/6x^5)(-1/5x)? I just cannot get it? the fifth power and negative signs are throwing me for a loop!

OpenStudy (anonymous):

1/30x^6.

jimthompson5910 (jim_thompson5910):

Multiply the coefficients -1/6 and -1/5 to get (-1/6)(-1/5)=1/30. This is the coefficient of the final answer. Now add the exponents: 5+1 = 6. This is the exponent for x of the final answer. So \[\large (-\frac{1}{6}x^5)(-\frac{1}{5}x)=\frac{1}{30}x^6\]

OpenStudy (anonymous):

Negative times negative = positive Multiply across. 6 * 5 = 30. x^5*x = x^6 =1/30x^6

OpenStudy (anonymous):

Why would you all write an answer if I answered it already?